### 06 Dec 2023 v1

@@ 3.4 Design of the loss function
4:     $w^{1}=\mathop{\arg\max}\limits_{{Q_{i}-\{v_{i}^{*},w_{i}^{0}\}}}\left(d(v_{i}^{*},{Q_{i}-\{v_{i}^{*},w_{i}^{0}\}})\right)$;
5:     $s=\frac{d(w^{0},w^{1})}{S}$;
6:     $\hat{w}_{i}^{\{1,...,S\}}=\{\Gamma_{(w^{0},w^{1})}(s\cdot 1),\Gamma_{(w^{0},w^{1})}(s\cdot 2),...,\Gamma_{(w^{0},w^{1})}(s\cdot S)\}$;
7:     $w_{i}^{*}=\mathop{\arg\min}\limits_{\hat{w}_{i}^{k}}\sum_{j=1}^{M_{i}}(d(z_{i}^{j},\Gamma_{(v_{i}^{*},\hat{w}_{i}^{k})}(\cdot)))^{2},k\in\{1,...,S\}$;
8:     Assign $w_{i}^{*}$ to $v_{i}^{*}$;
9:  end while
10:  Calculate $d(v_{i}^{*},w_{i}^{*})$ with Formula(2);
11:  for $k=1$ to $K$ do
12:     Sample $\hat{s}_{i}^{k}\sim U(0,d(v_{i}^{*},w_{i}^{*}))$;
13:     $\hat{z}_{i}^{k}=\cos(\hat{s}_{i}^{k})\cdot v_{i}^{*}+\sin(\hat{s}_{i}^{k})\frac{w_{i}^{*}-v_{i}^{*}\cdot\cos\theta_{(v_{i}^{*},w_{i}^{*})}}{\sin\theta_{(v_{i}^{*},w_{i}^{*})}}$;
14:  end for
@@ 4.1 Efficiently performing GCFA
CoMatch 62.37±0.12 68.10±0.21 80.71±0.13 82.84±0.24 88.24±0.05 90.95±0.17 93.64±0.11 94.39±0.21
$\Pi$-Model 54.04±0.23 69.29±0.34 82.90±0.16 84.12±0.19 80.66±0.13 82.56±0.22 92.90±0.14 92.77±0.37
Mean Teacher 70.50±0.22 75.96±0.14 81.95±0.11 83.68±0.26 81.21±0.17 84.33±0.19 92.64±0.03 92.43±0.25
Pseudo Labeling 73.51±0.25 76.29±0.22 82.61±0.15 84.42±0.21 81.92±0.21 82.44±0.26 94.15±0.12 94.57±0.23
MixMatch 73.87±0.32 76.06±0.15 84.07±0.21 84.58±0.34 81.81±0.25 85.07±0.19 93.83±0.16 94.31±0.31
VAT 77.33±0.17 80.40±0.32 83.91±0.07 85.43±0.31 89.92±0.04 90.21±0.11 95.28±0.05 95.14±0.36
Dash 78.74±0.23 82.69±0.23 86.46±0.03 87.17±0.16 93.20±0.05 93.55±0.21 95.30±0.07 95.90±0.25
CRMatch 79.84±0.15 80.54±0.26 86.12±0.17 87.16±0.22 92.79±0.08 93.31±0.32 95.41±0.13 95.89±0.29
UDA 80.51±0.14 82.52±0.31 86.31±0.15 87.50±0.18 91.74±0.11 92.13±0.25 95.33±0.12 95.78±0.22
FixMatch 80.59±0.25 82.96±0.23 86.76±0.04 87.46±0.36 91.40±0.12 91.84±0.23 95.08±0.09 95.35±0.35
FlexMatch 80.78±0.15 82.86±0.25 86.12±0.15 87.16±0.19 92.57±0.21 93.08±0.15 95.75±0.05 95.63±0.27
AdaMatch 81.93±0.17 83.78±0.22 85.84±0.23 87.13±0.32 90.95±0.12 92.96±0.23 95.12±0.03 95.58±0.32
FreeMatch 82.22±0.13 83.34±0.25 86.27±0.23 87.56±0.28 93.57±0.15 94.07±0.24 95.73±0.04 96.11±0.12
SimMatch 82.48±0.12 83.57±0.34 86.98±0.06 87.88±0.37 92.78±0.18 93.18±0.16 95.49±0.11 95.77±0.24
SoftMatch 82.72±0.11 83.11±0.28 86.83±0.17 87.81±0.26 93.27±0.11 93.68±0.12 95.72±0.16 95.55±0.34
@@ 4.3 The influence of machine learning models on GCFA
KNN 79.54 91.11 96.29 96.75
SVC 89.02 91.29 96.72 96.85
DecisionTree 56.76 58.51 85.62 86.03
ExtraTree 51.37 66.07 78.68 85.89
RandomForest 91.36 91.23 96.09 96.37
Bagging 74.52 77.87 94.38 94.52
GradientBoost 71.32 76.41 88.48 89.81
MLP 90.00 93.31 96.45 97.12

### 10 Dec 2023 v2

@@ 3.4 Design of the loss function
4:     $\hat{w}_{i}^{\{1,...,S\}}=\{\Gamma_{(w^{0},w^{1})}(s\cdot 1),$ $\qquad\Gamma_{(w^{0},w^{1})}(s\cdot 2),...,\Gamma_{(w^{0},w^{1})}(s\cdot S)\}$;
5:     $w_{i}^{*}=\mathop{\arg\min}\limits_{\hat{w}_{i}^{k}}\sum_{j=1}^{M_{i}}(d(z_{i}^{j},\Gamma_{(v_{i}^{*},\hat{w}_{i}^{k})}(\cdot)))^{2},$ $\qquad k\in\{1,...,S\}$;
6:     Assign $w_{i}^{*}$ to $v_{i}^{*}$;
7:  end while
8:  Calculate $d(v_{i}^{*},w_{i}^{*})$ with Formula(2);
9:  for $k=1$ to $K$ do
10:     Sample $\hat{s}_{i}^{k}\sim U(0,d(v_{i}^{*},w_{i}^{*}))$;
11:     $\hat{z}_{i}^{k}=\cos(\hat{s}_{i}^{k})\cdot v_{i}^{*}+\sin(\hat{s}_{i}^{k})\frac{w_{i}^{*}-v_{i}^{*}\cdot\cos\theta_{(v_{i}^{*},w_{i}^{*})}}{\sin\theta_{(v_{i}^{*},w_{i}^{*})}}$;
12:  end for
@@ 4.1 Efficiently performing GCFA
CoMatch[46] 62.37±0.12 68.10±0.21 80.71±0.13 82.84±0.24 88.24±0.05 90.95±0.17 93.64±0.11 94.39±0.21
$\Pi$-Model[47] 54.04±0.23 69.29±0.34 82.90±0.16 84.12±0.19 80.66±0.13 82.56±0.22 92.90±0.14 92.77±0.37
Mean Teacher[48] 70.50±0.22 75.96±0.14 81.95±0.11 83.68±0.26 81.21±0.17 84.33±0.19 92.64±0.03 92.43±0.25
Pseudo Labeling[49] 73.51±0.25 76.29±0.22 82.61±0.15 84.42±0.21 81.92±0.21 82.44±0.26 94.15±0.12 94.57±0.23
MixMatch[50] 73.87±0.32 76.06±0.15 84.07±0.21 84.58±0.34 81.81±0.25 85.07±0.19 93.83±0.16 94.31±0.31
VAT[51] 77.33±0.17 80.40±0.32 83.91±0.07 85.43±0.31 89.92±0.04 90.21±0.11 95.28±0.05 95.14±0.36
Dash[52] 78.74±0.23 82.69±0.23 86.46±0.03 87.17±0.16 93.20±0.05 93.55±0.21 95.30±0.07 95.90±0.25
CRMatch[53] 79.84±0.15 80.54±0.26 86.12±0.17 87.16±0.22 92.79±0.08 93.31±0.32 95.41±0.13 95.89±0.29
UDA[54] 80.51±0.14 82.52±0.31 86.31±0.15 87.50±0.18 91.74±0.11 92.13±0.25 95.33±0.12 95.78±0.22
FixMatch[55] 80.59±0.25 82.96±0.23 86.76±0.04 87.46±0.36 91.40±0.12 91.84±0.23 95.08±0.09 95.35±0.35
FlexMatch[56] 80.78±0.15 82.86±0.25 86.12±0.15 87.16±0.19 92.57±0.21 93.08±0.15 95.75±0.05 95.63±0.27
AdaMatch[58] 81.93±0.17 83.78±0.22 85.84±0.23 87.13±0.32 90.95±0.12 92.96±0.23 95.12±0.03 95.58±0.32
FreeMatch[57] 82.22±0.13 83.34±0.25 86.27±0.23 87.56±0.28 93.57±0.15 94.07±0.24 95.73±0.04 96.11±0.12
SimMatch[59] 82.48±0.12 83.57±0.34 86.98±0.06 87.88±0.37 92.78±0.18 93.18±0.16 95.49±0.11 95.77±0.24
SoftMatch[60] 82.72±0.11 83.11±0.28 86.83±0.17 87.81±0.26 93.27±0.11 93.68±0.12 95.72±0.16 95.55±0.34
@@ 4.3 The influence of machine learning models on GCFA
KNN[65] 79.54 91.11 96.29 96.75
SVC[66] 89.02 91.29 96.72 96.85
DecisionTree[67] 56.76 58.51 85.62 86.03
ExtraTree[68] 51.37 66.07 78.68 85.89
RandomForest[69] 91.36 91.23 96.09 96.37
Bagging[70] 74.52 77.87 94.38 94.52
GradientBoost[71] 71.32 76.41 88.48 89.81
MLP[72] 90.00 93.31 96.45 97.12